INAF queue observing, 2016 Feb 28-29

Observer: R. Carini, M. Faccini
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn
Instrument: LBC, LUCI

Summary

At the sunset the sky is clear but the clouds are coming. Wind at the limit (5m/s)

Night duration (12\xB0 el twilight) 19:30-5:30 ( 02:30-12:30 UT) = 10 h 00m

Observing time: 8 h (80%)

Weather time loss: 1 h (10%)

technical time loss 1 h (10%)

Data Summary

Project_RUNSorted ascending Objects Filter ExpTime Notes
bias LBC      
Castellano MACS J017 LBC-b(U_spec); LBC-r( V) 0h 30 m the last ob done is the 7
deBarros M103GLASS2c 200 H+K grating, zJspec 1h 10 m the last image taken during twilight
Flat LBC U_spec, V, B; V, I   just the U_spec images are good
flat and arc LUCI        
Marconi IZw18 LBC-b(V); LBC-r(I) 10m 6th ob done
Phot std SA95 LBC-b(U_spec, V,B), LBC-r( V,r,i,I)
Raimondo UmaII LBC-b(U_spec, B); LBC-r( r) 1h 56 m the last ob done is the 5
telluric HIP59861 200 H+K grating, zJspec    

Details

01:30 start Flat Fields

02:05 photometric standard SA95
  • dofpia
  • 02:10 send the ob
  • seeing 0.85''

2:23 slew to Marconi's program.
  • send copointing IZw18_copointing.ob
  • dofpia
  • 02:30 in the middle of the dofpia the machine obs2 is stucked, we pass to obs3
  • the AS controls that everything is ok and continue the dofpia
  • 02:40 lbcrangebal
  • 02:46 run science IZw18_6.ob
  • seeing 0.7''
  • 02:53 some cirrus start forming

03:00 we slew to Raimondo's program
  • run UMaII _f3_UBr_point.ob
  • dofpia, lbrangebal
  • 03:23 UMaII _f3_UBr.ob
  • 03:32 seeing 0.7'' on LBC-red
  • 03:40 seeing 0.62'' on LBC-red
  • 03:56 seeing 0.7''-0.8'' on LBC-red

04:03 run dofpia on UMaII _f4_UBr.ob
  • 04:09 run UMaII _f4_UBr.ob
  • 04:20 seeing 0.73'' on LBC-red

04:59 run dofpia on UMaII _f5_UBr.ob
  • 05:05 run UMaII _f5_UBr.ob
  • 05:23 seeing 0.65'' on LBC-red
  • 05:34 seeing 0.63'' on LBC-red
  • 05:40 some clouds start forming

05:55 we slew to Castellano A program MACS J0717.5
  • run copointing wt10_318_point.ob
  • dofpia, lbrangebal
  • 06:05 run Castellano_MACSJ0717_UV_7.ob
  • 06:22 seeing 0.75'''
  • 06:39 seeing 0.76''
  • 06:47 dofpia
  • 06:51 run Castellano_MACSJ0717_UV_8.ob
  • 07:00 dofpia again, the stars in the red channel are a little bit elongate (eccentricity is about 0.8-0.11) (the last dofpia failed in the red channel)
  • 07:06 run Castellano_MACSJ0717_UV_8.ob
  • 07:11 seeing 1.0''
  • 07:17 the clouds are increaseing
  • 07:22 the counts of the backgroud are increaseing due to the combination of the effects of moon and clouds , the images in the red are not good but in the blue are also good, we wait the acquisition of the last image in the blue and we stop
  • 07:30 stop

07:30 slew to LUCI, deBarros' program
  • 07:55 the TO send the preset
  • there are a loto of clouds, the sky is pretty bad.

8:09 load the Telluric script and slewing to the field. GO M103GLASS2c _tell
  • 8:15 abort collimation due to impossibility to collimate in the cloudy sky ( clouds moving fast from west )
  • the sky is covered, problem with acquisition and resolved (luci2.20162229.0194-0197), lost guiding star...but we have a spectrum of the telluric star luci2.20162229.0198, and after few seconds the scripts stopped...
  • 08:39 the sky is clearer, the telescope is guided, we continue the script luci2.0199-0201
  • seeing is about 0.8''
  • 08:43 offset failed, images from 0202 to 0205 are not good
  • 08:48 go again the script for the telluric but the sky is covered again and the telescope can not guide. We can not observe

09:09 the sky is a little bit clearer, we try again the acquisition of the telluric
  • the TO searches a pointing star
  • load the script for the telluric
  • 09:15 acquisition of the telluric star luci2.0206-0208 but there are problems with the focus
  • the AS trys to initialize the detector focus but the problems still remain

9:38 we decide to take the scientific data by hand luci2.0209-0214 we forget the dithering
  • we follow the dithering of the PI luci2.0215-0217 are ok
  • luci2.0218-0217 are not good, there are not the spectrum

9:50 stop the technical problem with automatic acquisition
  • with the AS we control the offset given by the PI luci2.0222-0226
  • we understand that the initial slit for the acquisition is wrong

10:20 acquisition telluric again luci2.0227-0229
  • 10:30 science in automatic mode luci2.0300-0307

10:41 slew to scientific target GO M103GLASS2c.xml
  • acquisition luci2.0238-0.241
  • 11:23 spectra luci2.0242-0244 but we do not see the spectra (to 0244)
  • load again the script and take the acquisition luci2.0245-0248
  • 11:40 again clouds but the seeing is very good 0.5'' luci2.0249-0254
  • run again the science GO M103GLASS2c.xml =luci2.0255-0260= seeing 0.5''-0.6''
  • 12:14 GO M103GLASS2c.xml luci2.0261-0262 the last one take during the twilight

12:59 begins the twilight, stop the observation

13:04 close the dome

13:28 begin flat for the program deBarros
  • flat without lamp luci2.0263-0267 with texp=30 sec
  • 13:32 flat with the lamp HALO3 luci2.0268 texp=30 sec, counts are low (about 1000)
  • take again the flats with HALO 3 and texp=60 sec luci2.0269-0273

13:41 LBC bias

13:45 flat without lamp and texp=60sec luci2.0274-0278
  • arc-neon luci2.0279-0283 texp=60sec
  • arc-argon =luci2.0284-0288 texp=5sec
  • flat without lamp and texp=5 sec ==luci2.0289-0293

Article text.

-- %USERSIG{RobertaCarini - 2016-02-29}%

Comments

 
Topic revision: r1 - 31 Jul 2024, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback